FREE RADICAL SPECTROSCOPY AND DYNAMICS BY LASER ABSORPTION TECHNIQUES

\underline{TREVOR J. SEARS}, CHEMISTRY DEPARTMENT, BROOKHAVEN NATIONAL LABORA-TORY, UPTON, NY 11973-5000.

Tunable continuous wave laser-based spectrometers have been used to record new spectra of molecular free radicals in the infrared and near infrared. This talk will concentrate on spectra of ethyl (C_{2}H_{5}) close to 530 cm^{−1}and a number of carbenes and other species at near infrared wavelengths. In C_{2}H_{5}, the spectrum is due to the CH_{2} out-of-plane rocking fundamental with complicated torsion-rotational structure that has now been analyzed to provide estimates for the torsional barrier and its change on vibrational excitation. At near-IR wavelengths, frequency modulation techniques can be used to increase the absorption sensitivity since the dominant noise source here is due to the probe laser; detector noise is negligible. This is not the case in the infrared where detector and laser source noise are comparable in magnitude. Examples from recent work on HCBr both at room temperature and in a jet-cooled source and simple metal containing radical species formed in an ablation source combined with a free jet expansion will be described. Acknowledgement: This work was carried out at Brookhaven National Laboratory under contract DE-AC02-98CH10886 with the U. S. Department of Energy and supported by its Division of Chemical Sciences, Office of Basic Energy Sciences.